Presently, Bill is 13 and his dad is 35. What will their ages be when the ration of their ages is 2/3?
Answer by jorel1380(3719)

13+p/35+p=2/3
3(13+p)=2(35+p)
39+3p=70+2p
p=70-39=31
Their ages will be 2/3 in 31 years..